# ZIP: Scalable Crowd Counting via Zero‑Inflated Poisson Modeling

### v2

@@ Abstract
Most crowd counting methods directly regress blockwise density maps using Mean Squared Error (MSE) losses. This practice has two key limitations: 1) it fails to account for the extreme spatial sparsity of annotations—over 95% of $8\times 8$ blocks are empty across standard benchmarks, so supervision signals in informative regions are diluted by the predominant zeros; 2) MSE corresponds to a Gaussian error model that poorly matches discrete, non‑negative count data. To address these issues, we introduce ZIP, a scalable crowd counting framework that models blockwise counts with a Zero‑Inflated Poisson likelihood: a zero‑inflation term learns the probability a block is structurally empty (handling excess zeros), while the Poisson component captures expected counts when people are present (respecting discreteness). We provide a generalization analysis showing a tighter risk bound for ZIP than MSE‑based losses and DMCount provided that the training resolution is moderately large. To assess the scalability of ZIP, we instantiate it on backbones spanning over 100×100$\times$100 × in parameters/compute. Experiments on ShanghaiTech A & B, UCF-QNRF, and NWPU-Crowd demonstrate that ZIP consistently surpasses state‑of‑the‑art methods across all model scales.
@@ Introduction
Crowd counting aims to estimate the number of people present in an image or a video frame. It underpins a broad range of real-world applications, including public safety and crowd management (Valencia et al. 2021; Saleh, Suandi, and Ibrahim 2015), urban planning (Waclawek et al. 2024), and intelligent transportation systems (McCarthy et al. 2025).
@@
Most modern crowd counting methods (Ma et al. 2019; Cheng et al. 2022; Han et al. 2023; Ranasinghe et al. 2024) construct blockwise training targets by first convolving point annotations with Gaussian kernels to create a pixel‑level density map, and then summing density values within each $B\times B$ block (e.g., $8\times 8$) to obtain blockwise density maps. Models are trained to regress these blockwise targets (usually with Mean Squared Error (MSE)) in lieu of raw pixel supervision to mitigate extreme pixel‑level sparsity.
@@
In these methods, however, the aggregated blockwise density maps remain highly sparse: across ShanghaiTech A & B (Zhang et al. 2016), UCF‑QNRF (Idrees et al. 2018), and NWPU‑Crowd (Wang et al. 2020c), over 95% of $8\times 8$ blocks contain zero people (empirical measurement). This zero dominance skews squared‑error losses, dilutes supervision from populated regions, and biases models toward under‑counting. These methods also suffer label noise introduced by Gaussian smoothing. Point annotations are spatially uncertain: a human click marks an approximate head center, and nearby pixels could be equally valid. Spreading each point with a local kernel is intended to make supervision tolerant to such subjective variation. However, since the true head extents are unknown, mismatched kernel bandwidths yield over‑ or under‑counts after block aggregation and thus inject noise into learning targets.
(a) Input Image.
(b) GT Density Map.
(c) Structural Zero Map.
(d) Predicted Density Map.
Figure 1: Visualization of ZIP on a ShanghaiTech B test image. Color scale (all panels): red = higher value, blue = lower value (per-panel normalization). In Fig. 1(c), red regions have high probability of being structurally empty (background, torso, peripheral head parts), blue indicates low structural-zero probability (candidate head centers). In Fig. 1(d), red regions correspond to high expected local crowd density near head centers; blue denotes near-zero density.
@@
Prior work has tried to mitigate sparsity by first segmenting foreground crowd regions and then counting only within those areas (Rong and Li 2021; Modolo et al. 2021; Guo et al. 2024). Such a practice introduces additional multi-task loss balancing complexity. Also, since crowd counting datasets are not annotated with segmentation masks, these approaches usually derive pseudo masks from Gaussian-smoothed density maps, inheriting the same head-size uncertainty and label noise discussed above.
@@
Figure 2: Illustration of structural and sampling zeros (best viewed in color). The image (synthesized for illustration purposes) is overlaid with a grid where each cell represents a spatial block. The red dot marks a ground-truth head annotation, and the block containing it is labeled with the red number “1”. Yellow zeros, which indicate sampling zeros, are assigned to blocks that correspond to the head-center region but receive zero count due to the point-based annotation protocol. Green zeros denote structural zeros, corresponding to background regions, non-head body parts or outer regions of the head that are not associated with any annotations.
@@
Additionally, MSE-based loss functions present a modeling mismatch. Blockwise targets represent discrete, non‑negative, and zero-heavy count data, whereas MSE implicitly assumes Gaussian residuals on continuous values. A Poisson model captures both the discreteness and non-negativity of counts, yet a standard Poisson struggles with the overwhelming presence of zeros. These gaps motivate a formulation that explicitly separates structural emptiness from stochastic counts, leading to our Zero-Inflated Poisson (ZIP) framework.
Our ZIP framework bypasses Gaussian smoothing and model per-block counts $y\in\mathbb{N}$ via a zero‑inflated Poisson:
$P(y=k|\pi,\lambda)=\begin{cases}\pi+(1-\pi)e^{-\lambda},\quad&k=0\\ (1-\pi)\frac{e^{-\lambda}\lambda^{k}}{k!},\quad&k>0\end{cases}$ (1)
with $\pi$, $\lambda$ predicted per block. In this way, we hypothesize that zeros in blockwise density maps arise from two distinct mechanisms (See Fig. 2 for illustration):
  • Structural zeros (modeled by $\pi$), which are blocks that are deterministically zero due to their semantic irrelevance. These include background (sky, buildings, etc.), body parts (limbs, chest, etc.) and peripheral head regions that do not correspond to head centers, and they comprise the majority of zero blocks.
  • Sampling zeros (modeled by $(1-\pi)e^{-\lambda}$): each head is annotated by a single point that we assign to exactly one supervision block. Neighboring blocks that also correspond to the head center therefore receive zero count. Modeling these annotation effects in the Poisson component absorbs the positional ambiguity that Gaussian smoothing was meant to solve, removing kernel bandwidth tuning.
Because structural and sampling zeros are not annotated separately, our ZIP model learns to distinguish these two types of zeros implicitly by optimizing the negative log-likelihood of the ZIP distribution, with $\lambda$ additionally supervised by a cross entropy loss to stabilize training. Empirically, learned zero‑inflation correlates with non-head-center areas. As illustrated in Fig. 1, the structural-zero map in 1(c) predicts nearly all non-head pixels as zeros, while preserving only the compact blue blobs centered on heads. Compared with segmentation-based methods, our ZIP-based method offers several notable advantages: 1) it does not rely on segmentation masks, thereby avoiding noise introduced by Gaussian smoothing; 2) it captures structural sparsity without assuming the presence of clear object boundaries, making it better suited for highly occluded and densely crowded scenes; 3) it eliminates the need to balance counting and segmentation, simplifying the training objective.
@@
Figure 3: Scalability of the proposed ZIP framework. The $x$-axis represents the number of model parameters (in millions), and the $y$-axis indicates the Mean Absolute Error (MAE; lower is better) on the validation set of the ShanghaiTech B dataset (Zhang et al. 2016). Each circle represents a specific model, where the radius corresponds to the number of floating point operations (FLOPs) required to process a single $1920\times 1080$ image during inference. ZIP models (in blue) exhibit a favorable trade-off between model size and accuracy, demonstrating better scalability compared to other lightweight (yellow) and heavyweight (green) models.
@@
Through theoretical risk bound analysis (See Theorem 1) and comprehensive empirical evaluation, we show that ZIP consistently outperforms Gaussian-smoothed MSE-based losses and DMCount (Wang et al. 2020a). To further assess the scalability of ZIP, we evaluate its performance with backbones of varying computational complexities, including MobileNet, MobileCLIP , and CLIP-ConvNeXt. As illustrated in Fig. 3, our ZIP framework consistently outperforms existing models in both accuracy (measured by MAE) and computational efficiency (measured by FLOPs) across different parameter scales, demonstrating its superior scalability. To summarize, our contributions are as follows:
@@
  • We conduct a theoretical analysis and experiments on four crowd counting benchmarks to validate the effectiveness of our ZIP framework. Results show that our base model, ZIP-Base, consistently outperforms existing state-of-the-art methods across four different datasets.
@@
Figure 4: Overview of the proposed ZIP framework. Given an input image $\boldsymbol{X}$ of spatial size $H\times W$, a backbone network extracts feature maps, which are shared by two parallel branches. In the Poisson branch (top), the features are passed to a $\boldsymbol{\lambda}$-head with a softmax activation, producing the blockwise count distribution $\boldsymbol{P}^{*}_{\boldsymbol{\lambda}}$ over $(n-1)$ positive bins. The Poisson rate per block is then computed as a weighted average of bin centers, yielding the final $\boldsymbol{\lambda}$ map. In the Zero Inflation branch (bottom), the same features are fed into a $\boldsymbol{\pi}$-head with a sigmoid activation to estimate the structural zero probability map $\boldsymbol{\pi}$ of spatial size $h\times w$. The predicted density map $\boldsymbol{Y}^{*}$ is defined as the expected value of the zero-inflated Poisson distribution: $\boldsymbol{Y}^{*}=(1-\boldsymbol{\pi})\otimes\boldsymbol{\lambda}$, where $\otimes$ denotes elementwise multiplication.
@@
In this work, we propose to incorporate Zero-Inflated Poisson (ZIP) regression to address the extreme sparsity of ground-truth density maps. Fig. 4 illustrates an overview of the proposed ZIP framework.
@@
(a) MAE
(b) RMSE
(c) NAE (%)
Figure 5: Performance of the VGG19-based structure on ShanghaiTech B under varying values of $\omega$ in the loss Eq. (11). Results indicate that $\omega=0.75$ yields the best MAE (6.34) and RMSE (9.79), while $\omega=1.00$ gives the lowest NAE (5.00%).
@@
$\boldsymbol{f}=\mathcal{F}(\boldsymbol{X})\in\mathbb{R}^{C\times h\times w}$ (2)
where $h=H//r$, $w=W//r$ denote the spatial size of the output.
Then, the feature map $\boldsymbol{f}$ is passed through a $\boldsymbol{\lambda}$-head $\mathcal{H}_{\boldsymbol{\lambda}}$ (implemented as a $1\times 1$ convolution), followed by a softmax activation to generate a probability distribution over the $n{-}1$ positive count bins:
$\boldsymbol{P}^{*}_{\boldsymbol{\lambda}}=\mathrm{Softmax}\left(\mathcal{H}_{\boldsymbol{\lambda}}(\boldsymbol{f})\right)\in\mathbb{R}^{(n-1)\times h\times w}.$ (3)
Note that we exclude the $\{0\}$ bin here, since the Poisson distribution assumes a strictly positive rate $\lambda>0$. We then compute the estimated Poisson rate map $\boldsymbol{\lambda}\in\mathbb{R}_{+}^{1\times h\times w}$ as the expected value of the bin centers:
@@
The learning of $\boldsymbol{P}^{*}_{\boldsymbol{\lambda}}$ is supervised by a cross-entropy loss over the positive regions of the ground-truth density map. Specifically, we define the set of positive-valued blocks as $\boldsymbol{Y}_{+}\coloneqq\boldsymbol{Y}[\boldsymbol{Y}>0]$, which arise from the Poisson component and are thus only related to $\boldsymbol{\lambda}$. Their corresponding probabilistic predictions $\boldsymbol{P}^{*}_{\boldsymbol{\lambda}+}$ can be obtained via $\boldsymbol{P}^{*}_{\boldsymbol{\lambda}+}=\boldsymbol{P}^{*}_{\boldsymbol{\lambda}}[:,\boldsymbol{Y}>0]$ The cross-entropy loss is calculated based on $\boldsymbol{P}^{*}_{\boldsymbol{\lambda}+}$ and $\boldsymbol{Y}_{+}$:
@@
In parallel, we introduce a $\boldsymbol{\pi}$-head, $\mathcal{H}_{\boldsymbol{\pi}}$, also implemented as a $1\times 1$ convolution, followed by a sigmoid activation to produce the structural zero probability map:
$\boldsymbol{\pi}=\sigma\left(\mathcal{H}_{\boldsymbol{\pi}}(\boldsymbol{f})\right)\in\mathbb{R}^{1\times h\times w}$ (6)
@@
$\mathcal{L}_{\mathrm{NLL}}=-\sum_{i=1}^{h}\sum_{j=1}^{w}\log P(\boldsymbol{Y}_{i,j}\mid\boldsymbol{\pi}_{i,j},\boldsymbol{\lambda}_{i,j})$ (7)
@@
It is important to note that we do not supervise the $\boldsymbol{\pi}$ and $\boldsymbol{\lambda}$ branches separately. Since structural and sampling zeros are not explicitly annotated in the ground-truth, we treat them as latent factors and optimize both branches jointly via the ZIP negative log-likelihood loss Eq. (7). This formulation allows the model to implicitly learn to disentangle the two types of zeros based on the statistical patterns in the data, without requiring additional annotations.
We use the expectation map $\boldsymbol{Y}^{*}=\mathbb{E}[\boldsymbol{V}|\boldsymbol{\pi},\boldsymbol{\lambda}]$ as the predicted density map, given by
$\boldsymbol{Y}^{*}_{i,j}=\mathbb{E}[\boldsymbol{V}_{i,j}\mid\boldsymbol{\pi}_{i,j},\boldsymbol{\lambda}_{i,j}]=(1-\boldsymbol{\pi})\otimes\boldsymbol{\lambda},$ (8)
@@
We define the count loss as the mean absolute error (MAE) between the predicted and ground-truth total counts:
@@
ZIP Has a Tighter Risk Bound
Theorem 1.
Assume the per-block counts follow a zero-inflated Poisson law with $\boldsymbol{\pi}>0$. Let $\mathcal{D}$ be the unknown ground-truth data distribution and $\mathcal{S}$ be the i.i.d. $K$ samples. Denote the hypothesis class of scalar regressors by $\mathcal{H}$ and the family of vector blockwise density predictors by =×$\mathcal{F}=\mathcal{H}^{\times hw}$. Let $f_{\mathcal{S}}^{\mathrm{NLL}}$ denotes an empirical risk minimizer of $\mathcal{L}_{\mathrm{NLL}}$ on $\mathcal{S}$, and $f_{\mathcal{D}}$ is the population minimizer (“Bayes optimal predictor”) for the same loss. Then, for any $\delta\in(0,1)$, with probability at least $1-\delta$ over the choice of the $K$ i.i.d. training examples $\mathcal{S}$, every $f\in\mathcal{F}$ satisfies
$\mathcal{R}_{\mathcal{D}}(f_{\mathcal{S}}^{\mathrm{NLL}},\mathcal{L}_{\mathrm{NLL}})-\mathcal{R}_{\mathcal{D}}(f_{\mathcal{D}},\mathcal{L}_{\mathrm{NLL}})$
$\leq 2\cdot hw\cdot C\cdot R_{\mathcal{S}}(\mathcal{H})+\mathcal{O}(\sqrt{1/K}),$ (12)
where $hw$ is the total number of blocks, $C$ is a dataset-specific constant, and $R_{\mathcal{S}}(\mathcal{H})$ is the empirical Rademacher complexity of the hypothesis class $\mathcal{H}$ on the sample set $\mathcal{S}$.
$\mathcal{R}_{\mathcal{D}}(f_{\mathcal{S}}^{\mathrm{OT}},\mathcal{L}_{\mathrm{OT}})-\mathcal{R}_{\mathcal{D}}(f_{\mathcal{D}},\mathcal{L}_{\mathrm{OT}})$
Hence, when $h$ or $w$ grows, the upper bound of ZIP (Eq. (12)) becomes tighter than that of OT (Eq. (13)). The proof of Theorem 1 can be found in the supplementary material.
Experiments
Variants. To evaluate the scalability and adaptability of our proposed framework, we construct five model variants of ZIP with different backbones, including MobileNetV4 (Qin et al. 2024), MobileCLIP (Vasu et al. 2024) and OpenCLIP’s ConvNeXt-Base (Liu et al. 2022; Cherti et al. 2023) These variants are suitable for different application scenarios, ranging from mobile-friendly to high-performance settings. The specifications are summarized in Table 1. All variants share the same head design and differ only in backbone complexity. This modularity allows ZIP to scale across computational budgets.
Setup. Our training configuration largely follows. We first preprocess all datasets so that the minimum edge length is no less than 448. We limit the maximum edge length of UCF-QNRF to be 2048 and that of NWPU-Crowd to be 3072. We apply a combination of data augmentation techniques, including random resized cropping, horizontal flipping, color jittering, and salt-and-pepper noise injection. For ShanghaiTech A and B, which consist of lower-resolution images, the crop size is set to $448\times 448$. For the higher-resolution datasets UCF-QNRF and NWPU-Crowd, we use a larger crop size of $672\times 672$. All models are optimized using the Adam optimizer with a weight decay of $41e-$ 4. The learning rate is linearly warmed up from $51e-$ 5 to $41e-$ 4 over the first 25 epochs, followed by cosine annealing with parameters $T_{0}=5$ and $T_{\mathrm{mult}}=2$.
Evaluation metrics. Following standard practice, we evaluate each model variant on all datasets using three commonly adopted metrics: Mean Absolute Error (MAE), Root Mean Squared Error (RMSE), and Normalized Absolute Error (NAE). The NAE metric ensures fair evaluation across both sparse and dense scenes by normalizing the absolute error with respect to the ground-truth count. These metrics are defined as follows:
$\mathrm{MAE}=\frac{1}{M}\sum_{m=1}^{M}\left|c_{m}-c_{m}^{*}\right|,$ (14)
$\mathrm{RMSE}=\sqrt{\frac{1}{M}\sum_{m=1}^{M}\left|c_{m}-c_{m}^{*}\right|^{2}},$ (15)
and
$\mathrm{NAE}=\frac{1}{\|\mathcal{M}_{+}\|}\sum_{m\in\mathcal{M}_{+}}\frac{\left|c_{m}-c_{m}^{*}\right|}{c_{m}},$ (16)
where $M$ denotes the total number of test images, $c_{m}$ is the ground-truth count for image $m$, and $c_{m}^{*}$ is the predicted count, computed via Eq. (9). The set $\mathcal{M}+\subseteq{1,2,\dots,M}$ includes only the indices for which $c_{m}>0$, and $\|\mathcal{M}_{+}\|$ denotes its cardinality.
@@
Table 1: ZIP variants with corresponding backbones, parameter sizes, and computational complexity measured on $1920\times 1080$ resolution.
Component MAE RMSE NAE
MAE 6.69 10.45 5.52%
DMCount (Wang et al. 2020a) 6.62 9.83 5.30%
Poisson NLL 6.57 9.86 5.47%
$\mathcal{L}_{\mathrm{NLL}}$ (ours) 6.36 10.32 5.00%
Table 2: Comparison of $\mathcal{L}_{\mathrm{NLL}}$ (Eq. (7)) with alternative choices in the total loss function (Eq. (11)) on ShanghaiTech B, while $\mathcal{L}_{\mathrm{CE}}$ and $\mathcal{L}_{\mathrm{count}}$ remain unchanged. Results show that our ZIP NLL achieves the lowest MAE and NAE.

### v3

@@ Abstract
Most crowd counting methods directly regress blockwise density maps using Mean Squared Error (MSE) losses. This practice has two key limitations: 1) it fails to account for the extreme spatial sparsity of annotations—over 95% of standard (8×8)(8$\times$ 8)( 8 × 8 ) blocks are empty across most benchmarks, so supervision signals in informative regions are diluted by the predominant zeros; 2) MSE corresponds to a Gaussian error model that poorly matches discrete, non‑negative count data. To address these issues, we introduce ZIP, a scalable crowd counting framework that models blockwise counts with a Zero‑Inflated Poisson likelihood: a zero‑inflation term learns the probability a block is structurally empty (handling excess zeros), while the Poisson component captures expected counts when people are present (respecting discreteness). We provide a generalization analysis showing a tighter risk bound for ZIP compared to that of MSE‑based losses and optimal-transport-based losses, provided that the training resolution is moderately large. To assess the scalability of ZIP, we apply it to backbones spanning over 100×100$\times$100 × in parameters/compute. Experiments on ShanghaiTech A & B, UCF-QNRF, and NWPU-Crowd demonstrate that ZIP consistently achieves state‑of‑the‑art performance across all compared model scales. Particularly, on UCF-QNRF, ZIP outperforms mPrompt by almost 3 MAE and 12 RMSE, demonstrating its effectiveness even in extremely dense crowd scenes.
(a) Input Image
(b) Ground Truth Density Map
(c) Structural Zero Map
(d) Predicted Density Map
Figure 1: The structural zero map output by ZIP can accurately segment non-head-central regions, thereby highlighting only candidate head-center areas. This visualization uses an image from the ShanghaiTech B test split. Color scale (all panels): red = higher value, blue = lower value (per-panel normalization). In (c), red regions have high probability of being structurally empty (background, torso, peripheral head parts), blue indicates low structural-zero probability (i.e., candidate head centers). In (d), red regions correspond to high expected local crowd density near head centers; blue denotes near-zero density. This structural-zero modeling addresses density map sparsity by masking out semantically empty regions.
@@ Introduction
Crowd counting aims to estimate the number of people present in an image or a video frame. It has a broad range of real-world applications, including public safety and crowd management (Valencia et al. 2021), and intelligent transportation systems (McCarthy et al. 2025). The majority of existing crowd counting methods (Ma et al. 2019; Cheng et al. 2022; Han et al. 2023; Ranasinghe et al. 2024) construct blockwise training targets by first convolving point annotations with Gaussian kernels to create a pixel‑level density map, and then summing density values within each $\times 8$ 8 block to obtain blockwise density maps. Models are trained to regress these blockwise targets (usually with Mean Squared Error (MSE)) in lieu of raw pixel supervision to mitigate extreme pixel‑level sparsity.
@@
In the above methods, however, the aggregated blockwise density maps remain highly sparse: across ShanghaiTech A & B (Zhang et al. 2016), UCF‑QNRF (Idrees et al. 2018), and NWPU‑Crowd (Wang et al. 2020c), over 95% of $\times 8$ 8 blocks contain zero people (empirical measurement). This zero dominance skews squared‑error losses, dilutes supervision from populated regions, and biases models toward under‑counting. These methods also suffer from label noise introduced by Gaussian smoothing. Point annotations are spatially uncertain: a human click marks an approximate head center, and nearby pixels could be equally valid. Spreading each point with a local kernel is intended to make supervision tolerant to such subjective variation. However, since the true head extents are unknown, mismatched kernel bandwidths yield over‑ or under‑counts after block aggregation and thus inject noise into learning targets.
@@
Prior work attempts to mitigate sparsity by first segmenting foreground crowd regions and then counting only within those areas (Rong and Li 2021; Modolo et al. 2021; Guo et al. 2024). Such a practice introduces additional multi-task loss balancing complexity. Also, since crowd counting datasets are not annotated with segmentation masks, these approaches usually derive pseudo masks from Gaussian-smoothed density maps, inheriting the same head-size uncertainty and label noise discussed above.
@@
Figure 2: Illustration of the concept of structural and sampling zeros in ZIP (best viewed in color). The image (synthesized for illustration purposes) is overlaid with a grid where each cell represents a spatial block. The red dot marks a ground-truth head annotation, and the block containing it is labeled with the red number “1”. Yellow zeros, which indicate sampling zeros, are assigned to blocks that correspond to the head-center region but receive zero count due to the point-based annotation protocol. Green zeros denote structural zeros, corresponding to background regions, non-head body parts or outer regions of the head that are not associated with any annotations.
@@
In addition, MSE-based loss functions present a modeling mismatch. Blockwise targets represent discrete, non‑negative, and zero-heavy count data, whereas MSE implicitly assumes Gaussian residuals on continuous values. A Poisson model can capture both the discreteness and non-negativity of counts, yet a standard Poisson struggles with the overwhelming presence of zeros. These gaps motivate a formulation that explicitly separates structural emptiness from stochastic counts, leading to our Zero-Inflated Poisson (ZIP) framework.
@@
Figure 3: Scalability of the proposed ZIP framework on ShanghaiTech B. Each circle represents a specific model, where the radius corresponds to the number of floating point operations (FLOPs) required to process a single 1920×10801920$\times$ 10801920 × 1080 image during inference. ZIP models (in blue) exhibit a favorable trade-off between model size and accuracy, demonstrating better scalability compared to other lightweight (yellow) and heavyweight (green) models.
Our ZIP framework hypothesizes that zeros in blockwise density maps arise from two distinct mechanisms :
  • Structural zeros, which are blocks that are deterministically zero due to their semantic irrelevance. These include background (sky, buildings, etc.), body parts (limbs, chest, etc.) and peripheral head regions that do not correspond to head centers, and they comprise the majority of zero blocks. Fig. 1(c) illustrates this type of zeros predicted by ZIP on ShanghaiTech B.
  • Sampling zeros: each head is annotated by a single point which is assigned to a unique supervision block. Neighboring blocks that also correspond to the head center therefore receive zero count. Modeling these annotation effects in the Poisson component absorbs the positional ambiguity that Gaussian smoothing is meant to solve, removing kernel bandwidth tuning.
Fig. 2 provides a schematic comparison of these two types of zeros. Because structural and sampling zeros are not annotated separately, our ZIP model learns to distinguish these two types of zeros implicitly by optimizing the negative log-likelihood of the ZIP distribution, with the Poisson rate additionally supervised by a cross entropy loss to stabilize training. Empirically, learned zero‑inflation correlates with non-head-center areas. As illustrated in Fig. 1 , the structural-zero map in 1(c) predicts nearly all non-head pixels as zeros, while preserving only the compact blue blobs centered on heads. Compared with segmentation-based methods, our ZIP-based method offers several notable advantages: 1) it does not rely on segmentation masks, thereby avoiding noise introduced by Gaussian smoothing; 2) it captures structural sparsity without assuming the presence of clear object boundaries, making it better suited for highly occluded and densely crowded scenes; 3) it eliminates the need to balance counting and segmentation, simplifying the training objective.
@@
Through theoretical risk bound analysis (see Theorem 1) and comprehensive empirical evaluation, we show that ZIP consistently outperforms Gaussian-smoothed MSE-based losses and DMCount (Wang et al. 2020a). To further assess the scalability of ZIP, we evaluate its performance with backbones of varying computational complexities, including MobileNet, MobileCLIP, and CLIP-ConvNeXt. As illustrated in Fig. 3, our ZIP framework consistently outperforms existing models in both accuracy (measured by Mean Absolute Error (MAE)) and computational efficiency (measured by FLOPs) across different parameter scales, demonstrating its superior scalability. To summarize, our contributions are:
@@
  • We present both theoretical risk bound analysis and thorough experiments on four crowd counting benchmarks to validate the effectiveness of our ZIP framework. Results show that our base model, ZIP-Base, consistently outperforms existing state-of-the-art methods across four different datasets.
@@
In this work, we propose to incorporate Zero-Inflated Poisson (ZIP) regression to address the extreme sparsity of ground-truth density maps. Our ZIP framework bypasses Gaussian smoothing and models per-block counts $y\in\mathbb{N}$ via a zero‑inflated Poisson :
$P(y=k|\pi,\lambda)=\begin{cases}\pi+(1-\pi)e^{-\lambda},\quad&k=0\\ (1-\pi)\cfrac{e^{-\lambda}\lambda^{k}}{k!},\quad&k>0\end{cases}$ (1)
with $\pi$, $\lambda$ predicted per block. Fig. 4 illustrates an overview of the proposed ZIP framework.
@@
$\boldsymbol{f}=\boldsymbol{F}(\boldsymbol{X})\in\mathbb{R}^{C\times h\times w}$ (2)
where $h=H//r$, $w=W//r$ denote the spatial size of the output, and $r$ represents the block size.
Then, the feature map $\boldsymbol{f}$ is passed through a $\boldsymbol{\lambda}$-head $\boldsymbol{}{H}_{\boldsymbol{\lambda}}$ (implemented as a $\times 1$ 1 convolution), followed by a softmax activation to generate a probability distribution over the $n{-}1$ positive count bins:
$\boldsymbol{P}^{*}_{\boldsymbol{\lambda}}=\mathrm{Softmax}\left(\boldsymbol{H}_{\boldsymbol{\lambda}}(\boldsymbol{f})\right)\in\mathbb{R}^{(n-1)\times h\times w}.$ (3)
Note that we exclude the $\mathcal{B}_{1}=\{0\}$ bin here, since the Poisson distribution assumes a strictly positive rate $\lambda>0$. We then compute the estimated Poisson rate map $\boldsymbol{\lambda}\in\mathbb{R}_{+}^{1\times h\times w}$ as the expected value of the bin centers:
@@
The learning of $\boldsymbol{P}^{*}_{\boldsymbol{\lambda}}$ is supervised by a cross-entropy loss over the positive regions of the ground-truth density map. Specifically, we define the set of positive-valued blocks as $\boldsymbol{Y}_{+}\coloneqq\boldsymbol{Y}[\boldsymbol{Y}>0]$, which arise from the Poisson component and are thus only related to $\boldsymbol{\lambda}$. Their corresponding probabilistic predictions $\boldsymbol{P}^{*}_{\boldsymbol{\lambda}+}$ can be obtained via $\boldsymbol{P}^{*}_{\boldsymbol{\lambda}+}=\boldsymbol{P}^{*}_{\boldsymbol{\lambda}}[:,\boldsymbol{Y}>0]$. The cross-entropy loss is calculated based on $\boldsymbol{P}^{*}_{\boldsymbol{\lambda}+}$ and $\boldsymbol{Y}_{+}$:
@@
In parallel, we introduce a $\boldsymbol{\pi}$-head, $\boldsymbol{H}_{\boldsymbol{\pi}}$, also implemented as a $\times 1$ 1 convolution, followed by a sigmoid activation to produce the structural zero probability map:
$\boldsymbol{\pi}=\sigma\left(\boldsymbol{H}_{\boldsymbol{\pi}}(\boldsymbol{f})\right)\in\mathbb{R}^{1\times h\times w}$ (6)
@@
$\mathcal{L}_{\mathrm{NLL}}=-\frac{1}{hw}\sum_{i=1}^{h}\sum_{j=1}^{w}\log P(\boldsymbol{Y}_{i,j}\mid\boldsymbol{\pi}_{i,j},\boldsymbol{\lambda}_{i,j})$ (7)
@@
It is important to note that we do not supervise the $\boldsymbol{\pi}$ and $\boldsymbol{\lambda}$ branches separately. Since structural and sampling zeros are not explicitly annotated in the ground-truth, we treat them as latent factors and optimize both branches jointly via the ZIP negative log-likelihood loss in Eq. (7). This formulation allows the model to implicitly learn to disentangle the two types of zeros based on the statistical patterns in the data, without requiring additional annotations.
We use the expectation map $\boldsymbol{Y}^{*}=\mathbb{E}[\boldsymbol{Y}|\boldsymbol{\pi},\boldsymbol{\lambda}]$ as the predicted density map, given by
$\boldsymbol{Y}^{*}=(1-\boldsymbol{\pi})\otimes\boldsymbol{\lambda}\in\mathbb{R}^{1\times h\times w}$ (8)
@@
We define the count loss as the MAE between the predicted and ground-truth total counts:
@@
Table 1: ZIP variants with corresponding backbones, parameter sizes, and computational complexity measured on a 1920×10801920$\times$ 10801920 × 1080 resolution.
ZIP has a Tighter Risk Bound
Theorem 1.
Assume that the per-block counts follow a zero-inflated Poisson (ZIP) distribution parameterized by blockwise parameters $\boldsymbol{\theta}=(\boldsymbol{\pi},\boldsymbol{\lambda})$, with $\boldsymbol{\pi}>0$. Let $\mathcal{H}$ denote the hypothesis class of scalar regressors (e.g., the shared backbone and prediction heads), and define the full hypothesis class as the blockwise product ×$\mathcal{F}\coloneqq\mathcal{H}^{\times hw}$, where $hw$ is the total number of spatial blocks. Let $f_{\mathcal{S}}^{\mathrm{NLL}}$ be the empirical risk minimizer of the ZIP negative log-likelihood loss $\mathcal{L}_{\mathrm{NLL}}$ over the sample $\mathcal{S}$, and let $f_{\mathcal{D}}^{\mathrm{NLL}}$ be the population minimizer of the same loss. Then, for any $\delta\in(0,1)$, with probability at least $1-\delta$ over the random draw of $\mathcal{S}$, the following generalization bound holds:
$\mathcal{R}_{\mathcal{D}}(f_{\mathcal{S}}^{\mathrm{NLL}},\mathcal{L}_{\mathrm{NLL}})-\mathcal{R}_{\mathcal{D}}(f_{\mathcal{D}}^{\mathrm{NLL}},\mathcal{L}_{\mathrm{NLL}})$
$\leq 2\cdot hw\cdot L\cdot R_{\mathcal{S}}(\mathcal{H})+\mathcal{O}(\sqrt{1/K}),$ (12)
where $L$ is the Lipschitz constant of the full-image ZIP loss (per Lemma 2 in the supplementary material), and $\mathcal{R}_{\mathcal{S}}(\mathcal{H})$ is the empirical Rademacher complexity of the base hypothesis class $\mathcal{H}$.
$\mathcal{R}_{\mathcal{D}}(f_{\mathcal{S}}^{\mathrm{OT}},\mathcal{L}_{\mathrm{OT}})-\mathcal{R}_{\mathcal{D}}(f_{\mathcal{D}}^{\mathrm{OT}},\mathcal{L}_{\mathrm{OT}})$
where $C_{\infty}$ is the maximum cost in the cost matrix in OT. Hence, when $h$ or $w$ grows, the upper bound of ZIP (Eq. (12)) becomes tighter than that of OT (Eq. (13)). The proof of Theorem 1 and the risk bound of the overall loss function Eq. (11) can be found in the supplementary material.
Experiments
Variants. To evaluate the scalability and adaptability of our proposed framework, we construct five model variants of ZIP with different backbones, including MobileNetV4 (Qin et al. 2024), MobileCLIP (Vasu et al. 2024) and OpenCLIP’s ConvNeXt-Base (Liu et al. 2022; Cherti et al. 2023). These variants are suitable for different application scenarios, ranging from mobile-friendly to high-performance settings. The specifications are summarized in Table 1. All variants share the same head design and differ only in backbone complexity. This modularity allows ZIP to scale across computational budgets.
[…]
Evaluation metrics. Following standard practice, we evaluate each model variant on all datasets using three commonly adopted metrics: MAE, RMSE, and Normalized Absolute Error (NAE) . The NAE metric ensures fair evaluation across both sparse and dense scenes by normalizing the absolute error with respect to the ground-truth count.
Figure 5: Performance of the VGG19-based structure on ShanghaiTech B under varying values of $\omega$ in Eq. (11). Results indicate that $\omega=1.00$ gives the lowest MAE (6.03).
Loss Function Component MAE RMSE NAE
MAE 6.39 10.66 5.34%
MSE 6.90 11.41 5.30%
DMCount (Wang et al. 2020a) 6.62 11.35 5.30%
Poisson NLL 6.21 10.25 5.02%
ZIP NLL (ours) 6.03 9.95 4.99%
Table 2: Ablation study on the proposed ZIP NLL loss in Eq. (7). We replace it with alternative choices in the total loss function in Eq. (11) on ShanghaiTech B, while $\mathcal{L}_{\mathrm{CE}}$ and $\mathcal{L}_{\mathrm{count}}$ remain unchanged. Results show that our ZIP NLL achieves the lowest MAE, RMSE, and NAE.